OpenStudy (anonymous):

What is the value of the expression 3–4? A. –81 B. –12 C. – D.

OpenStudy (michele_laino):

hint: \[\Large {3^{ - 4}} = \frac{1}{{{3^4}}} = \frac{1}{{3 \cdot 3 \cdot 3 \cdot 3}} = ...?\]

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@Michele_Laino nvm can i have the answer

OpenStudy (michele_laino):

\[\Large \begin{gathered} {3^{ - 4}} = \frac{1}{{{3^4}}} = \frac{1}{{3 \cdot 3 \cdot 3 \cdot 3}} = \frac{1}{{\left( {3 \cdot 3} \right) \cdot \left( {3 \cdot 3} \right)}} \hfill \\ \hfill \\ = \frac{1}{{9 \cdot 9}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

what is 9*9=...?

OpenStudy (anonymous):

81

OpenStudy (michele_laino):

correct! so your answer is: \[\Large \begin{gathered} {3^{ - 4}} = \frac{1}{{{3^4}}} = \frac{1}{{3 \cdot 3 \cdot 3 \cdot 3}} = \frac{1}{{\left( {3 \cdot 3} \right) \cdot \left( {3 \cdot 3} \right)}} \hfill \\ \hfill \\ = \frac{1}{{9 \cdot 9}} = \frac{1}{{81}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

my answer is 1/8?

OpenStudy (michele_laino):

your answer is: \[\Large \frac{1}{{81}}\]

OpenStudy (anonymous):

k thx